# An Efficient Difference-of-Convex Solver for Privacy Funnel

### v1

@@ IV Evaluation
(a) DCA (14) $q=1$ versus Submodular [3]
(b) DCA (14) $q=1$ versus (13) $q=2$ (ridge)
Figure 1: Comparing the characterization of the privacy-utility trade-off between the proposed DCA solvers and the state-of-the-art method [3]. Under known joint distribution of $P(X,Y)$ settings, given by (25).
@@ IV-A Known Distribution
Observe that the $y_{0},y_{1}$ of this distribution are easy to infer if observing $x_{0},x_{1}$ whereas there is ambiguity in inferring $y_{0},y_{2}$ if $x_{2}$ is observed instead.
The goal of our evaluation is to characterize the privacy-utility trade-off by plotting the lowest achieved $I(Z;Y)$ for an obtained $I(Z;X)$. The characterized trade-off is displayed as the information plane [1, 4], which depends on both $P(X,Y)$ and the algorithm used. We evaluate our DCA solver (6a) and compare it to the state-of-the-art [1, 3] when $P(X,Y)$ is known. These baselines only consider deterministic $P(z|x)=\boldsymbol{1}\{x\in z\}$, that is, $z$ categorizes $X$. Also, they are greedy algorithms that merge two or more $z\in\mathcal{Z}$ each step for a lower PF Lagrangian. Contrary to them, our solver is non-greedy with feasible solutions covers the (smoothed) probability simplice. For these baselines, it suffices to compare to the solver in [3] since the solver in [1] is a special case of it.
In the proposed solver (6a), the range of cardinality $|\mathcal{Z}|$ follows the Carathéodory Theorem, which is should satisfy $|\mathcal{Z}|\leq\max\{|\mathcal{X}|,|\mathcal{Y}|\}+1$ [24]. For each cardinality, we initialize $P(Z|X)$ by randomly sample a uniformly distributed source, followed by normalization to obtain a starting point. Then, we fix a trade-off parameter $\beta\in[0.1,10.0]$ and a regularization coefficient $\alpha\in[0.1,10.0]$, where each set consists of $16$ geometrically spaced points. For each set of hyperparameters, we run our solver until either the consecutive loss values satisfy $|\mathcal{L}^{k}-\mathcal{L}^{k+1}|\leq 10^{-6}$ or a maximum number of iterations $M=10^{4}$ is reached. This procedure is repeated for $10$ times. We compute the obtained metrics $I(Z;X),I(Z;Y)$ offline.
In Fig. 1a, our solver (6a) characterizes the trade-off better than the baseline solvers since we not only cover their solutions but achieve more non-trivial points that are infeasible to them. In Fig. 1b we compare the two proposed DCA solvers with different implementation of the inner problem, i.e., (13) with $q=2$ (ridge) versus (14) with $q=1$. While the ridge approach performs worse than the other solver in the characterization of the trade-off, it obtains a better trade-off point $I(Z;X)\approx 1$ bit and $I(Z;Y)\approx 0.3$ bits. Moreover, the ridge solver benefits from the optimized off-the-shelf solver [19], hence converges faster in our empirical evaluation. It is worth noting that both approaches outperform [3] in characterizing the trade-off under different regularization criterion, which is the strength of our DCA approach.
@@ IV-B Unknown Distribution
Figure 2: Privacy-utility trade-off in unknown joint distribution settings. The proposed DCA solver is compared to the CPF baseline [7].
For unknown $P(X,Y)$, we consider cases where empirical samples of the joint distribution are available (datasets) $\mathcal{D}_{N}:=\{(x^{(n)},y^{(n)})|(x^{(n)},y^{(n)})\sim P(X,Y)\}_{n=1}^{N}$. Our DCA solver is compared to the state-of-the-art Conditional Privacy Funnel (CPF) [7] baseline.
@@
we evaluate (18a) on two datasets: MNIST [25] and Fashion MNIST [26]. Both datasets consist of gray-scale images of $28\times 28$ pixels. Each pixel is normalized to the range $[0,1]$. Also, both datasets have labeled $60000$ training and $10000$ testing samples uniformly distributed between $10$ classes respectively. The data format of MNIST is images hand-written digits. Here, we let the digits $Y$ be the private information and the images are treated as the public information $X$. As for Fashion MNIST, it has $10$ categories of clothing. We refer to [26] for details. We set the categories as $Y$ and the gray-scale images as $X$.
The two datasets are applied to the scenario: a sender learns (supervised) an encoder-decoder pair with training dataset based on the PF objective. Then the sender encodes the $X$ of the testing set and the released compressed codes $Z$. A receiver observes the coded $Z$ and reconstruct $\hat{X}$ with $P_{\phi}(X|Z)$, obtained from a secure channel before the inference phase. Meanwhile, an adversary also observes the codes but does not know $P_{\phi}(X|Z)$. The adversary therefore learns $\hat{Y}$ from clustering $Z$. Here, we assume the cardinality of $|Y|$ is known for simplicity. In short, the legitimate users’ goal is to reconstruct $\hat{X}$ from $Z$ whereas the adversary clusters $Z$ and obtains $\hat{Y}$ that should reveal $Y$.
In our evaluation, we measure the reconstruction quality (utility) by the Peak Signal-to-Noise power Ratio (PSNR), defined as:$PSNR:=10\log{X_{\max}/MSE}$, where $X_{\max}$ is the maximum possible pixel value ($1$ in our case) and $MSE:=\frac{1}{N}\sum_{i=1}^{N}\lVert\boldsymbol{x}_{i}-\hat{\boldsymbol{x}}\rVert^{2}$. As for the privacy leakage, the adversary uses UMAP [27] to project $Z$ to a $\mathbb{R}^{2}$ space, followed by a KMeans clustering [19]. Note that this two-step approach achieves the state-of-the-art clustering performance for both MNIST and Fashion-MNIST datasets [27]. Then we compute the adversary’s success rate by an offline label matching.
@@
The proposed and compared methods are implemented as deep neural networks. Since our focus is on comparing the claimed PF objective, the encoders and decoders are almost aligned for all compared methods. For both methods, we implement the encoder as a neural network with fully connected (FC) layers whose outputs correspond to that of the standard VAE [20] encoder. We have a separate set of parameters for learning a prior $P_{\theta}(Z)=\mathcal{N}(0,\boldsymbol{\nu}_{\theta})$. As for the decoder, another neural network with FC layers is the common reconstruction module for both approaches. For the differences, our method has a separate linear classifier whereas the CPF baseline expand the decoder’s input by concatenating code samples $\boldsymbol{z}$ with the one-hot representation of $\boldsymbol{e}_{y}$ ($\boldsymbol{e}_{i}$ is the $i^{th}$ elementary vector).
We train a model for $100$ epochs with a learning rate at $3\times 10^{-4}$ using a standard ADAM optimizer [28]. The mini-batch size is $256$. For the dimension of $Z$, we vary $d_{z}\in\{32,64,128,256\}$ for both methods. The hyperparameter range for our method is set to $[5\times 10^{-4},5\times 10^{-2}]$ for both $\alpha$ and $\beta$ in (17a), whereas it is $[10^{-3},10]$ for the CPF baseline. Each hyperparameter pair is trained for $3$ time with a different random seed. Then we report the average in the results. The experiment details are included in Appendix E. For reproduction, our source codes in available online111available at: https://github.com/hui811116/dcaPF-torch.
@@ Results
The results are shown in Fig. 2, where the highest adversary success rate of an achieved PSNR for each method is reported. Note that the results include all combinations of the hyperparameters and the set of $d_{z}$. Fig. 2a corresponds to the MNIST dataset whereas the results for Fashion-MNIST dataset is shown in Fig. 2b. Clearly, in both figures, our method outperforms the CPF baseline since the adversary has higher error rate in clustering $Z$ from our approach at a comparable reconstruction quality of $X$. In addition to better robustness, it is worth noting that in the inference phase, our approach is completely independent to the private information, in sharp contrast to the CPF baseline. These demonstrate the strength of the proposed DCA method. Finally, for extended evaluation and visualization of the compressed features, we refer to Appendix F and Appendix G for details.
@@ V Conclusions
We leverage the DC structure of the PF method and develop efficient solvers based on DCA. In known distribution settings, our solver resembles both ridge regression and sparse recovery. Empirically, our solver outperforms the state-of-the-art greedy solvers in characterizing the privacy-utility trade-off. As for the unknown distribution settings, the insights from the proposed DC approach allows compliance to the fundamental Markov relation, so that our approach applies to scenarios without labeled private information after training. The evaluation of the proposed solver of the MNIST and the Fashion-MNIST datasets demonstrates significantly improved privacy-utility trade-off than the compared methods. For future work, we empirically find that the compression ratio, defined as the ratio of the dimensionality of the release information to the public information, significantly affects the privacy-utility trade-off. A fundamental characterization in this direction will be the focus of a follow-up work.
@@ Appendix A Derivation for a known joint distribution
$\frac{\partial g(\boldsymbol{p}_{z|x})}{\partial\boldsymbol{p}(z|x)}:=P(x)\left\{\log{P(z)}+1+\beta\log{\frac{P(z|x)}{P(z)}}\right\}.$
@@
By definition, $\boldsymbol{B}_{y|x}$ is the matrix form of $P(Y|X)$ with $(i,j)$-entry $P(y_{i}|x_{j})$, so the above is equivalent to:
@@ Appendix D Proof of Theorem 1
The proof follows that in [30, Theorem 1]. Which is included here for completeness.
@@
$\nabla f(p^{k+1})-\nabla g(p^{k})=0.$ (27)
Denote the objective function as $\mathcal{L}(p)=f(p)-g(p)$ where $f,g$ are convex functions of $p$. We consider the following:
$\mathcal{L}(p^{k})-\mathcal{L}(p^{k+1})$
$=$ $f(p^{k})-g(p^{k})-f(p^{k+1})+g(p^{k+1})$
$\geq$ $\langle\nabla f(p^{k+1})-\nabla g(p^{k}),p^{k}-p^{k+1}\rangle,$
where the first inequality follows from the convexity of $f$ and $g$. Then by applying (27), we show that:
$\mathcal{L}(p^{k})-\mathcal{L}(p^{k+1})\geq 0.$ (28)
The above implies that the consecutive updates from (8) results in a non-increasing sequence of the objective values. For $k\geq K_{0}\in\mathbb{N}$ sufficiently large, we have a sequence of $\{p^{k}\}_{k\geq K_{0}}$ such that $\{\mathcal{L}(p^{k})\}_{k\geq K_{0}}$ converges to a $\mathcal{L}^{*}$ as $k\rightarrow\infty$. Denote the converged solutions $p^{*}\in\Omega^{*}$ where $\Omega^{*}:=\{p|\mathcal{L}(p)=\mathcal{L}^{*}\}$, then $\Omega^{*}$ is a set of stationary points whose objective value is $\mathcal{L}^{*}$. Now, because $\mathcal{L}$ is non-convex, $\Omega^{*}$ could include both local and global stationary points.

### v3

@@ IV Evaluation
(a) DCA (13) with $q=1$ versus Submodular [3]
(b) DCA (13) with $q=1$ versus (12) with $q=2$ (ridge)
Figure 1: Comparing the characterization of the privacy-utility trade-off between the proposed DCA solvers and the state-of-the-art method [3]. The joint distribution $P(X,Y)$ is known and is given by (18).
@@ IV-A Known Distribution
Note that in (18), $y_{0},y_{1}$ are easy to infer if observing $x_{0},x_{1}$ whereas there is ambiguity in inferring $y_{0},y_{2}$ if $x_{2}$ is observed instead.
The goal of our evaluation is to characterize the privacy-utility trade-off by plotting the lowest achieved $I(Z;Y)$ for an obtained $I(Z;X)$. The characterized trade-off is displayed on the information plane [1, 4], which depends on both $P(X,Y)$ and the algorithm used. We evaluate our DCA solver (5a) and compare it to the state-of-the-art solvers [1, 3]. These baselines only consider deterministic $P(z|x)=\boldsymbol{1}\{x\in z\}$, that is, $z$ categorizes $x,\forall x\in X$. These greedy algorithms merge two or more $z\in\mathcal{Z}$ iteratively for lower PF Lagrangian loss. Contrary to them, our solver is non-greedy with feasible solutions cover the (smoothed) probability simplice. For these baselines, it suffices to compare to [3] since [1] is a special case.
In the proposed solver (5a), the range of cardinality $|\mathcal{Z}|$ follows the Carathéodory Theorem, i.e., $|\mathcal{Z}|\leq\max\{|\mathcal{X}|,|\mathcal{Y}|\}+1$ [24]. For each cardinality, we initialize $P(Z|X)$ by randomly sample a uniformly distributed source in $[0,1]$, followed by normalization to obtain a starting point. Then, we fix a trade-off parameter $\beta\in[0.1,10.0]$ and a regularization coefficient $\alpha\in[0.1,10.0]$, where each set consists of $16$ geometrically spaced points. For each pair of hyperparameters, we run our solver until either the consecutive loss values satisfy $|\mathcal{L}^{k}-\mathcal{L}^{k+1}|\leq 10^{-6}$ or a maximum number of iterations $K=10^{4}$ is reached. This procedure is repeated for $10$ times. We compute the obtained metrics $I(Z;X),I(Z;Y)$ offline.
In Fig. 1a, our solver (5a) characterizes the trade-off better than the baseline solvers since we not only cover their solutions but achieve more non-trivial points that are infeasible to them. In Fig. 1b we compare the two proposed DCA solvers with different implementations, i.e., (12) with $q=2$ (ridge) versus (13) with $q=1$. While the ridge approach performs worse than the other solver in characterizing the trade-off, it obtains a better trade-off point at $I(Z;X)\approx 1$ bit and $I(Z;Y)\approx 0.3$ bits. Moreover, the ridge solver benefits from the optimized off-the-shelf solver [19], hence converges faster in our empirical evaluation. It is worth noting that both approaches outperform [3] in characterizing the trade-off under different regularization criterion, demonstrating the strength of our DCA approach.
@@ IV-B Unknown Distribution
Figure 2: Privacy-utility trade-off in unknown joint distribution settings. The proposed DCA solver is compared to the CPF [7] and DVPF [8] baselines.
For unknown $P(X,Y)$, we consider cases where empirical samples of the joint distribution are available (datasets) $\mathcal{D}_{N}:=\{(x^{(n)},y^{(n)})|(x^{(n)},y^{(n)})\sim P(X,Y)\}_{n=1}^{N}$. Our DCA solver is compared to the state-of-the-art Conditional Privacy Funnel (CPF) [7] and Deep Variational Privacy Funnel (DVPF) [8] baselines.
@@
we evaluate our solver (16a) on MNIST [25] and Fashion MNIST [26] datasets. Both datasets consist of gray-scale images of $28\times 28$ pixels. Each pixel is normalized to the range $[0,1]$. There are labeled $60000$ training and $10000$ testing samples uniformly distributed between $10$ classes for each dataset. Specifically, the MNIST dataset consists of images of hand-written digits. Here, we let the digits $Y$ be the private information and the images are treated as the public information $X$. As for Fashion MNIST, it has $10$ categories of clothing. We refer to [26] for details. We set the categories as $Y$ and the gray-scale images as $X$.
We consider the following scenario: a sender learns (supervised) an encoder-decoder pair with training dataset based on the PF objective. The decoder is shared to a receiver through a secure channel before inference phase. Then the sender encodes $X$ of the testing set as the released codes $Z$. A receiver observes the coded $Z$ and reconstruct $\hat{X}$ with the decoder $P_{\phi}(X|Z)$. Meanwhile, an adversary also observes the codes but does not know $P_{\phi}(X|Z)$. The adversary therefore learns $\hat{Y}$ from clustering $Z$. Here, we assume the cardinality of $|Y|$ is known for simplicity (relaxed with cluster identification algorithms [27]). The legitimate users’ goal is to reconstruct $\hat{X}$ from $Z$ whereas the adversary clusters $Z$ as $\hat{Y}$ that should reveal $Y$.
In our evaluation, we measure the reconstruction quality (utility) by the Peak Signal-to-Noise power Ratio (PSNR), defined as:$PSNR:=10\log{X_{\max}/MSE}$, where $X_{\max}$ is the maximum possible pixel value ($1$ in our case) and $MSE:=\frac{1}{N}\sum_{i=1}^{N}\lVert\boldsymbol{x}_{i}-\hat{\boldsymbol{x}}\rVert_{2}^{2}$. As for the privacy leakage, the adversary uses UMAP [28] to project $Z$ to a $\mathbb{R}^{2}$ space, followed by KMeans clustering [19]. Note that this two-step approach achieves the state-of-the-art clustering performance for both MNIST and Fashion-MNIST datasets [28]. Then we compute the adversary’s success rate from offline label matching.
@@
The proposed and compared methods are implemented as deep neural networks. Since our focus is on comparing the PF objective, the encoders and decoders are carefully aligned for all compared methods. We implement the encoder as a neural network with fully connected (FC) layers as the standard VAE [20] encoder. We have a separate set of parameters for learning a priori $P_{\theta}(Z)=\mathcal{N}(\boldsymbol{0},\boldsymbol{\nu}_{\theta})$. As for the decoder, we employ another neural network with FC layers as the reconstruction module. For the differences, our method has a separate linear classifier whereas the CPF baseline expands the decoder’s input by concatenating code samples $\boldsymbol{z}$ with the one-hot representation of target variables, i.e., $\boldsymbol{e}_{y}$ ($\boldsymbol{e}_{i}$ is the $i^{th}$ elementary vector). For the DVPF baseline, extra adversarial generative networks are implemented according to [8].
We train each model for $100$ epochs with a learning rate at $3\times 10^{-4}$ using a standard ADAM optimizer [29]. The mini-batch size is $256$. For the dimension of $Z$, we vary $d_{z}\in\{32,64,128,256\}$ for all methods. The range of hyperparameters for our method is set to $[5\times 10^{-4},5\times 10^{-2}]$ for both $\alpha$ and $\beta$ in (16a), whereas it is $[10^{-3},10]$ for the CPF and DVPF baselines. Each hyperparameter pair is trained for $3$ times with a different random seed and we report the average in the results. The experiment details are included in Appendix E. For reproduction, our source codes in available online111available at: https://github.com/hui811116/dcaPF-torch.
@@ Results
In Fig. 2, we report the highest adversary success rate of an achieved PSNR for each method. Note that the results include all combinations of the hyperparameters and the set of $d_{z}$. Fig. 2a corresponds to the MNIST dataset whereas the results for Fashion-MNIST dataset is shown in Fig. 2b. In both figures, our method clearly outperforms the CPF and DVPF baselines since the adversary has lower success rate in clustering $Z$ from our approach at a comparable reconstruction quality of $X$. This demonstrates that our approach characterizes the privacy-utility trade-off better than both baselines. In addition to better robustness, our approach is more efficient in training phase than the DVPF baseline since we require no parameters for generative and discriminative modules whose optimization follows by a complex six-step algorithm. As for in inference phase, our approach is completely independent to the private information, in sharp contrast to the CPF baseline. For extended evaluation and visualization of the compressed features, we refer to Appendix F and Appendix G for details.
@@ V Conclusions
We leverage the DC structure of the PF method and develop efficient solvers based on DCA. In known distribution settings, our solver resembles both ridge regression and sparse recovery. Empirically, our solver outperforms the state-of-the-art greedy solvers in characterizing the privacy-utility trade-off. As for the unknown distribution settings, the insights from the proposed DC approach complies the fundamental Markov relation, so our approach applies to scenarios without labeled private information after training in contrast to compared methods. The evaluation of the proposed solver in MNIST and the Fashion-MNIST datasets demonstrates significantly improved characterization of the privacy-utility trade-off over the compared methods. For future work, we empirically find that the compression ratio ($|\mathcal{Z}|/|\mathcal{X}|$), significantly affects the characterization of the privacy-utility trade-off. A fundamental study in this direction will be the focus of a follow-up work.
@@ Appendix A Derivation for a known joint distribution
$\frac{\partial g(\boldsymbol{p}_{z|x})}{\partial P(z|x)}:=P(x)\left\{\log{P(z)}+1+\beta\log{\frac{P(z|x)}{P(z)}}\right\}.$
@@
By definition, $\boldsymbol{B}_{y|x}:=\boldsymbol{I}_{|\mathcal{Z}|}\otimes\boldsymbol{Q}_{y|x}$ where $\boldsymbol{Q}_{y|x}$ is the matrix form of $P(Y|X)$ with $(i,j)$-entry $P(y_{i}|x_{j})$, so the above is equivalent to:
@@ Appendix D Proof of Theorem 1
The proof follows that in [31, Theorem 1]. Which is included here for completeness.
First we establish the following lemma, using the fact that the linear equation $\boldsymbol{p}_{z}=\boldsymbol{A}_{x}\boldsymbol{p}_{z|x}$ represents $P(Z)=\sum_{x\in\mathcal{X}}P(X)P(Z|X)$, where $\boldsymbol{A}_{x}:=\boldsymbol{I}_{|\mathcal{Z}|}\otimes\boldsymbol{p}_{x}$ denotes the matrix form of the known marginal $P(X)$:
Lemma 1 (Restricted Convexity)
Let $g(\boldsymbol{p}_{z|x})$ be defined as in (5a), then we have:
$g(\boldsymbol{p}_{z|x})$$\geq$ $g(\boldsymbol{q}_{z|x})+\langle\nabla g(\boldsymbol{q}_{z|x}),\boldsymbol{p}_{z|x}-\boldsymbol{q}_{z|x}\rangle$
$+\frac{1}{2}\lVert\boldsymbol{A}_{x}(\boldsymbol{p}_{z|x}-\boldsymbol{q}_{z|x})\rVert^{2}_{2},$
where $\boldsymbol{p}_{z|x},\boldsymbol{q}_{z|x}\in\Omega_{z|x}$, with $\Omega_{z|x}$ denotes the feasible solution set of the conditional probability $P(Z|X)$ in vector form.
Proof:
It is well-known that the negative Shannon entropy $-H(X):=-H(\boldsymbol{p}_{x})$ is strongly convex with respect to the probability vector $\boldsymbol{p}_{x}$. Observe that $-H(Z)$ is a component of $g(\boldsymbol{p}_{z|x})$ and that the marginal $\boldsymbol{p}_{z}=\boldsymbol{A}_{x}\boldsymbol{p}_{z|x}$, and that the other components are also convex. The proof is complete. ∎
@@
$\nabla f(\boldsymbol{p}_{z|x}^{k+1})-\nabla g(\boldsymbol{p}_{z|x}^{k})=0.$ (20)
Denote the objective function as $\mathcal{L}(\boldsymbol{p})=f(\boldsymbol{p})-g(\boldsymbol{p})$ where $f,g$ are convex functions of $\boldsymbol{p}\in\Omega_{z|x}$. We consider the following:
$\mathcal{L}(\boldsymbol{p}_{z|x}^{k})-\mathcal{L}(\boldsymbol{p}_{z|x}^{k+1})$
$=$ $f(\boldsymbol{p}_{z|x}^{k})-g(\boldsymbol{p}_{z|x}^{k})-f(\boldsymbol{p}_{z|x}^{k+1})+g(\boldsymbol{p}_{z|x}^{k+1})$
$\geq$ $\langle\nabla f(\boldsymbol{p}_{z|x}^{k+1})-\nabla g(\boldsymbol{p}_{z|x}^{k}),\boldsymbol{p}_{z|x}^{k}-\boldsymbol{p}_{z|x}^{k+1}\rangle$
$+\frac{1}{2}\lVert\boldsymbol{A}_{x}(\boldsymbol{p}_{z|x}^{k}-\boldsymbol{p}_{z|x}^{k+1})\rVert^{2}_{2},$
where the first inequality follows from the convexity of $f$, followed by the restricted convexity of $g$ as shown in Lemma 1. Then by applying (20), we show that:
$\mathcal{L}(p^{k})-\mathcal{L}(p^{k+1})\geq\frac{1}{2}\lVert\boldsymbol{A}_{x}(\boldsymbol{p}^{k}_{z|x}-\boldsymbol{p}_{z|x}^{k+1})\rVert_{2}^{2}.$ (21)
The above implies that the consecutive updates from (7) results in a monotone decreasing sequence of the objective values. Summing both side from $k=0$ to $k=K-1$ sufficiently large, we have:
$\mathcal{L}(\boldsymbol{p}^{0}_{z|x})-\mathcal{L}(\boldsymbol{p}^{K}_{z|x})\geq\sum_{k=0}^{K-1}\lVert\boldsymbol{A}_{x}(\boldsymbol{p}^{k}_{z|x}-\boldsymbol{p}_{z|x}^{k+1})\rVert^{2}_{2}.$
For $K\rightarrow\infty$ sufficiently large, and under mild smoothness conditions [4], $\mathcal{L}$ is lower-semi-continuous, and hence the left hand side is finite, which in turns implies that the Cauchy sequence in the right hand side is bounded. Therefore, we have a sequence $\{\boldsymbol{A}_{x}\boldsymbol{p}_{z|x}^{k}\}_{k\geq K_{0}}$, obtained from the DCA algorithm (7) such that $\{\mathcal{L}(\boldsymbol{p}_{z|x}^{k})\}_{k\geq K_{0}}$ converges to a $\mathcal{L}^{*}$ as $k\rightarrow\infty$. Denote the converged solutions $\boldsymbol{A}_{x}\boldsymbol{p}_{z|x}^{*}\in\Omega^{*}$ where $\Omega^{*}:=\{\boldsymbol{A}_{x}\boldsymbol{q}|\nabla f(\boldsymbol{q})=\nabla g(\boldsymbol{q}),\boldsymbol{q}\in\Omega_{z|x}\}$, then $\Omega^{*}$ is a set of stationary points whose objective value is $\mathcal{L}^{*}$. Now, because $\mathcal{L}$ is non-convex, $\Omega^{*}$ could include both local and global stationary points.